x 0 1 2 3 y 8 1 −6 −13


What is an equation that describes the relationship?



1
Expert's answer
2022-01-10T12:58:14-0500

Let "y=kx+b"


"y(0)=b=8=>b=8"

"y=kx+8"

"y(1)=k+8=1=>k=-7"

Then "y=-7x+8."

Check


"y(2)=-7(2)+8=-6=-6, True"

"y(3)=-7(3)+8=-13=-13, True"

Answer:


"y=-7x+8"